# Efficient & Correct Predictive Equivalence for Decision Trees

### v2

@@ 7 Conclusions
Recent work by McTavish et al. (McTavish et al., 2025c, a) demonstrated the importance of eliminating predictive equivalent decision trees from the Rashomon set. Indeed, this work showed that a significant percentage of decision trees in the Rashomon sets for different classification problems are predictively equivalent. Specifically, the work of McTavish et al. proposed an approach (MBDSR) for solving a number of computational problems of interest for working with Rashomon sets of DTs, including deciding predictive equivalence of DTs. At its core, the MBDSR approach consists of using the well-known Quine-McCluskey (QM) method for computing minimimum-size DNF representations of DTs. The additional algorithms proposed by McTavish et al. run in linear time on the size of the computed minimimum-size DNF.
In this paper, we show two critical issues with the MBDSR approach proposed by McTavish et al. (McTavish et al., 2025c, a). As a first issue, we show that the MBDSR approach is unnecessarily complex in that it requires running the QM method for solving a problem that is hard for the second level of the polynomial hierarchy. In addition, we also prove that the worst-case exponential running time and space of the QM method can be exercised in the case of decision trees. In contrast, in this paper, we propose novel polynomial-time algorithms, in the size of the original DT, for all of the computational problems studied in the work of McTavish et al. (McTavish et al., 2025c, a). The experiments confirm negligible running times in practice for all the algorithms proposed in this paper. As a second issue, we prove that the MBDSR approach is flawed. Concretely, we demonstrate that the use of the QM method for deciding predictive equivalence can produce incorrect results. Therefore, the MBDSR is provably inefficient and, assuming that the set covering step of the QM method is not guaranteed to be deterministic, it may produce incorrect results; this depends on the implementation of the QM method that is used.
@@ Acknowledgments.
This work was supported in part by the Spanish Government under grant PID 2023-152814OB-I00, and by ICREA starting funds. The first author (JMS) also acknowledges the extra incentive provided by the ERC in not funding this research.

### v3

@@ 7 Conclusions
Recent work by McTavish et al. (McTavish et al., 2025c, a) demonstrated the importance of eliminating predictive equivalent decision trees from the Rashomon set. Indeed, this work showed that a significant percentage of decision trees in the Rashomon sets for different classification problems are predictively equivalent. Specifically, the work of McTavish et al. proposed an approach McTavish et al. (2025c, a) for solving a number of computational problems of interest for working with Rashomon sets of DTs, including deciding predictive equivalence of DTs. At its core, the work of McTavish et al. consists of using the well-known Quine-McCluskey (QM) method for computing minimimum-size DNF representations of DTs. The additional algorithms proposed by McTavish et al. run in linear time on the size of the computed minimimum-size DNF.
In this paper, we show two critical issues, one related with the efficiency of the algorithms proposed by McTavish et al. (McTavish et al., 2025c, a), and the other with the use of the QM method for predictive equivalence. As a first issue, we show that the approach McTavish et al. is unnecessarily complex in that it requires running the QM method for solving a problem that is hard for the second level of the polynomial hierarchy. In addition, we also prove that the worst-case exponential running time and space of the QM method can be exercised in the case of decision trees. In contrast, in this paper, we propose novel polynomial-time algorithms, in the size of the original DT, for all of the computational problems studied in the work of McTavish et al. (McTavish et al., 2025c, a). The experiments confirm negligible running times in practice for all the algorithms proposed in this paper. As a second issue, we argue that the use of the set-covering step of the QM method may produce incorrect results when deciding predictive equivalence. In their work, McTavish et al. address the issue of correctness by imposing constraints on how set-covering is implemented. However, one of the constraints required by McTavish et al. may be difficult to formally guarantee in practice; this will depend on a number of aspects that our paper details.
@@ Acknowledgments.
McTavish et al. provided detailed comments to earlier versions of this document. This work was supported in part by the Spanish Government under grant PID 2023-152814OB-I00, and by ICREA starting funds. The first author (JMS) also acknowledges the extra incentive provided by the ERC in not funding this research.